OpenStudy (anonymous):

Find the least common denominator for the following fractions: 3/4, 19/20, 1/6, 3/5, 8/15

OpenStudy (anonymous):

4 = 2^2 20 = 4 * 5 = 2^2 * 5 6 = 2 * 3 5 = 5 15 = 3 * 5 LCD = 2^2 * 3 * 5 = 60

OpenStudy (anonymous):

60 the answer

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

u welcome my cupcake

OpenStudy (anonymous):

50 Ahem, the site has rules. if you dont know already lets thank jigglypuff for letting you know... mmm if you're not familiar with the rules... http://openstudy.com/code-of-conduct This site has a policy for "Give Help, Not Answers" violations of such policies could potentially result in an account suspension or ban. Don't game the system! Keep medals and achievements valuable, so please avoid: questions specifically designed to inflate the medal count of the users who are participating.